Dynamics question:
61kg woman pulling a sled at a constant velocity by means of a rope at 37°. The tension on the rope is 120 N, and the mass of the sled is 55 kg.
a) find the normal force on the sled and on the woman
b) find the kinetic friction acting on the sled in the static friction acting on the woman
c) find the coefficient of kinetic friction for the sled
1
Expert's answer
2020-07-10T10:23:22-0400

a)


"N_s=m_sg-120\cdot\sin37°=55\cdot9.81-120\cdot\sin37°\approx467N";


"N_w=m_wg=61\cdot9.81\approx598N"


b)


"-F_{fr}+120\cdot \cos37°=0\to F_{fr}=120\cdot\cos37°\approx96N"


c)


"F_{fr}=\mu N_s\to\mu=\frac{F_{fr}}{N_s}=\frac{96}{467}=0.206"